A tour group split into two groups when waiting in line for food at a fast food counter. The first group brought 8 slices of pizza and 4 soft drinks for $36.68. The second group bought 6 slices of pizza and 6 soft drinks for $35.52. How much does one slice of pizza cost?
I assume the equations are 8x+4y=36.68 and 6x+6y=35.52.
Is the answer $3.25?
Thank you in advance!

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

x=pizza
8x+4y=36.68;-24x-12y=-110.04
6x+6y=35.52;12x+12y=71.04
The sum is -12x=-39
and x=$3.25 confirming.

Answer by MathTherapy(10809)


A tour group split into two groups when waiting in line for food at a fast food counter. The first group brought 8 slices of pizza and 4 soft drinks for $36.68. The second group bought 6 slices of pizza and 6 soft drinks for $35.52. How much does one slice of pizza cost?
I assume the equations are 8x+4y=36.68 and 6x+6y=35.52.
Is the answer $3.25?
Thank you in advance!
Yes, it is! Good job!
8x + 4y = 36.68______4(2x + y) = 4(9.17)_____2x + y = 9.17 ------ eq (i)
6x + 6y = 35.52______6(x + y) = 6(5.92)_______x + y = 5.92 ------ eq (ii)
x = $3.25------- Subtracting eq (ii) from eq (i)